Given the economy's marginal propensity to consume = .8

(Ceteris paribus) by how much will national income increase when the government raises the level of transfer payments by $100?

$100.

$500.

$80.

$400.
Answer #1

Ans: $500

Explanation:

Multiplier = 1 / ( 1 -MPC)

= 1 / ( 1 - 0.8 )

= 1 / 0.2

= 5

As government spending multiplier is 5 , the national income will increase by $500 when the government raises the level of transfer payments by $100.

Increase in national income = $100 * 5 = $500